OpenStudy (anonymous):

Using a particular magnifying glass, all length measurements appear to be six times their original sizes. What percent increase in length measurements is this?

OpenStudy (anonymous):

@DanJS

OpenStudy (anonymous):

@Owlcoffee

OpenStudy (anonymous):

@melodycrazysouthern

OpenStudy (anonymous):

@Zale101

OpenStudy (danjs):

it says 6 times

OpenStudy (danjs):

oh percent, meaning of 100, to get 6 from a fraction of 100 you need 600/100, that is 600%

OpenStudy (melodycrazysouthern):

? confusing

OpenStudy (anonymous):

600 is wrong

OpenStudy (danjs):

the lengths are 6 times larger, or 600% the original lengths

OpenStudy (anonymous):

600% is the wrong answer

OpenStudy (anonymous):

6 times the original size

OpenStudy (danjs):

hmm make an example, if an original length is 1 , and the zoomed length is 6 times that or 6 6 is what percent of 1 6 = (x/100) * 1 x=600

OpenStudy (anonymous):

thats weird.... its telling me 6,60, and 600 are all wrong answers @DanJS

OpenStudy (danjs):

did you type the % sign at the end

OpenStudy (danjs):

those computer things are not good

OpenStudy (anonymous):

its already on there for me, i just need to put a number in

OpenStudy (danjs):

not sure, if you are 1 ft tall, and appear to be 6 ft tall , that is 6 times the original length, or 6*100/100 = 600/100

OpenStudy (danjs):

600 per-cent

OpenStudy (anonymous):

yeah i don't know the answer... thats wrong

OpenStudy (anonymous):

oh well